OpenStudy (anonymous):

How to answer x+2 divided by x-2 = 4/8? What are the steps? Thanks.

hartnn (hartnn):

write 4/8 =1/2 then cross-multiply

OpenStudy (bahrom7893):

multiply both sides by 8(x-2). You'll get: 8(x+2) = 4(x-2)

OpenStudy (waleed_imtiaz):

means (x+2)/(x-2)=4/8...... Simply cross multiply and solve for x....

hartnn (hartnn):

u know how to cross-multiply ? if \(\large \frac{a}{b}=\frac{c}{d} \quad\text{then}\quad ad=bc \)

OpenStudy (anonymous):

So, if I cross multiply, is the answer x=4?

hartnn (hartnn):

sure?

OpenStudy (anonymous):

If the answer is x = 4, then I understand. Please confirm.

hartnn (hartnn):

2(x+2) = x-2 2x+4 =x-2 2x-x=-2-4 x=-6

OpenStudy (anonymous):

I will study this tonight. Thanks.

hartnn (hartnn):

welcome :)